OpenStudy (anonymous):

The length of a rectangle is 3cm longer than its width. If the perimeter of the rectangle is 30cm , find its area.

OpenStudy (anonymous):

lenght=x+3=y; width=x; 2x+2y=30; 2x+2x+6=30; x=6; then the area A=6*(6+3)=54 cm^2